OpenStudy (anonymous):

(6,5), (8,2), (x,11) = 15 Solve for x

OpenStudy (rob1525):

Are those points on a graph?

OpenStudy (anonymous):

yes

OpenStudy (rob1525):

I've never seen points set equal to a number. What math are you in?

OpenStudy (anonymous):

Algerbra 2, its asking for area of a triangle

OpenStudy (rob1525):

okay that makes mores sense, let me see.

OpenStudy (rob1525):

sorry its early its taking me a little longer. this is what ive got so far. if you find line for the points (6,5) (8,2) and then find the line perpendicular to one of the point youll get a triangle with a 90 degree angle. Now you can use the area formula 1/2bh to find area.

OpenStudy (rob1525):

Did you get it yet. I plotted a horizontal line on my graphing calc and included the line y=2/3x+1 and y=-3/2x+14. you can see that the horizontal line 11 crosses the line 2/3x+1 at point (15,11)